# DiffSHEG: A Diffusion-Based Approach for Real-Time Speech-driven Holistic 3D Expression and Gesture Generation

### 09 Jan 2024 v1

@@ 1 Introduction
In this work, we propose DiffSHEG, a unified Diffusion-based Holistic Speech-driven Expression and Gesture generation framework. To capture the joint distribution, DiffSHEG utilizes diffusion models [16] with a unified expression-gesture denoising network. As shown in Figure 2, our denoising network consists of two modules, an audio encoder, and a Transformer-based Uni-direction Expression-Gesture (UniEG) generator, which has a unidirectional flow from expression to gesture. The proposed framework ensures a natural temporal alignment between speech and motion, leveraging the relationship between expressions and gestures. Furthermore, we introduce a Fast Out-Painting-based Partial Autoregressive Sampling (FOPPAS) method to synthesize arbitrary long sequences efficiently. FOPPAS enables real-time streaming sequence generation without conditioning on previous frames during training, providing more flexibility and efficiency.
@@
Figure 2: DiffSHEG framework overview. Left: Audio Encoders and UniEG-Transformer Denoiser. Given an audio clip, we encode the audio into a low-level feature Mel-Spectrogram and a high-level HuBERT feature. The audio features are concatenated with other optional conditions, such as text, and then fed into the UniEG Transformer Denoiser. The denoising block fuses the conditions with noisy motion at diffusion step t and feeds it into style-aware transformers to get the predicted noises. The uni-directional condition flow is enforced from expression to gesture. Right: The detailed architecture of style-aware Transformer encoder and motion-condition fusion residual block.
@@ 2 Related Work
Co-speech Expression Generation. Co-speech Expression Generation, also known as talking head/face generation, has been a topic of interest in the field. Early related methods primarily relied on rule-based procedural approaches [10, 30, 41, 45], which required significant manual effort despite offering explicit and flexible control. With the advancements in deep learning, data-driven methods [7, 28, 21, 40, 32, 9, 20, 52, 35] have emerged, focusing on generating images that correspond to audio input. However, these methods often produce distorted pixels and inconsistent 3D results. On another front, some approaches [21, 34] propose animating 3D faces. More recently, FaceFormer [12] and CodeTalker [44] have utilized Transformer models to animate 3D facial vertices using designed attention maps for temporal alignment. While lip movement is closely tied to speech, most data-driven methods for expression animation design deterministic mappings, lacking diversity in eye movement and facial expressions. In contrast to these methods, our approach aims to leverage generative models to generate diverse facial expressions with precise lip movements driven by speech.
@@
Holistic Co-Speech Expression and Gesture Generation. There are some recent methods [14, 48] exploring the holistic joint generation of expression and gesture. Habibie et al. [14] first try to synthesize 3D facial and gesture motion at the same time. They propose a CNN-based framework that shares the same speech encoder and uses three decoders to generate distinct whole-body keypoints (i.e., for face, hand, body), followed by a discriminator to provide adversarial loss during training. However, this method learns deterministic mapping, which violates the nature of many-to-many mapping from speech to gesture motion. Yi et al. [48] use an encoder-decoder network with pre-trained Wav2vec [4] audio feature for expression generation and VQ-VAE for gesture generation. Nevertheless, They generate the expressions and gestures separately, losing the coherency between them. Moreover, the VQ-VAE-based methods have limited motion tokens, which may fail to generate motion as diverse and agile as real motions.
@@ 3 Method
How to model the expression-gesture joint distribution can be key to this task. A naïve way is directly concatenating the two vectors together and passing them to deep networks simultaneously, such that the features of both can be shared by each other. However, according to our experiments, this does not work well and leads to sub-optimal results. We hypothesize that the information flow from gesture to expression would interfere with the mapping from speech to expression. Intuitively, expressions can serve as cues for gestures since the gestures usually behave consistently with the emotions and speech inferred from expression [39] and lip movement [11]. On the contrary, gestures can hardly affect expressions, especially the lips, which have a strong and exclusive correlation with speech. Imagine when you express surprise, you may have widened eyes, opened mouth, and raised eyebrows, with frozen or startled gestures such as clasping hands and leaning backward. Therefore, we propose a unified framework based on diffusion models with uni-directional information flow from expression to gesture for unified co-speech expression and gesture generation to capture their joint distributions. At inference time, to generate arbitrary-long motion sequences, we propose an fast out-painting-based partial autoregressive sampling (FOPPAS) strategy, which generates partially overlapping clips one by one with a smooth motion at the intersection. Additionally, FOPPAS also features its real-time streaming inference ability. In the following sections, we will introduce the problem formulation and diffusion model [16], and then elaborate on the design of our proposed framework.
@@ 3.1 Problem Formulation
Given an arbitrary-long (streaming) audio, we aim to generate realistic, synchronous, and diverse gestures as well as expressions. To get training samples, we cut the aligned audio and motion sequences into clips with the same time duration using sliding windows. For each $N$-frame clip, we encode the corresponding audio clip to audio features $\mathbf{A}=[{\mathbf{a}_{1},\dots,\mathbf{a}_{N}}]$. Gesture clip $\mathbf{G}=[{\mathbf{g}_{1},\dots,\mathbf{g}_{N}}]$ is represented by joint rotations in axis-angle rotation format where $\mathbf{g}_{i}\in\mathbb{R}^{3J}$ and $J$ is the total joint number. Expression clip $\mathbf{E}=[{\mathbf{e}_{1},\dots,\mathbf{e}_{N}}]$ is represented by the blend shape weights where $\mathbf{e}_{i}\in\mathbb{R}^{C_{exp}}$ and $C_{exp}$ is the channel dimension of expression. The holistic motion sequence $\mathbf{M}=Concat(\mathbf{G},\mathbf{E})$, where each motion frame $\mathbf{m}_{i}\in\mathbb{R}^{3J+C_{exp}}$. The training objective of our framework is to reconstruct the motion clip $\mathbf{M}$ conditioned on audio clip $\mathbf{A}$, which is similar to the Multi-Input Multi-Output (MIMO) setting [33]. At inference time, we generate realistic, diverse, and well-aligned co-speech expressions and gestures with motion clips smoothly connected.
@@ 3.2 Preliminary: Diffusion Models
DiffSHEG uses diffusion models [16] for this task which consists of a diffusion and a denoising process. For explicitness, we substitute notations of expression, gesture, and motion clip $\mathbf{E},\mathbf{G},\mathbf{M}$ to $\mathbf{x}_{0}^{E},\mathbf{x}_{0}^{G},\mathbf{x}_{0}$. Given a motion clip distribution $p(\mathbf{x}_{0})\in\mathbb{R}^{N\times(3J+C_{exp})}$, our goal is to train a model parameterized by $\theta$ to approximate $p(\mathbf{x}_{0})$.
Diffusion Process. In the diffusion process, models progressively corrupts input data $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$ according to a predefined schedule $\beta_{t}\in(0,1)$, eventually turning data distribution into an isotropic Gaussian in $T$ steps. Each forward transition can be assumed as
@@
where the full forward process can be written as
@@
Denoising Process. In the denoising process, models learn to invert the noising procedure so that it can turn noise into real data distribution at inference. The corresponding backward process can be written as
@@
$\mathcal{L}_{t}$ $=\mathbb{E}_{\mathbf{x}_{0},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t\right)\right\|^{2}\right].$ (4)
@@ 3.3 DiffSHEG Framework
Motion-Speech Fusion Residual Block. In order to condition the motion on speech and align them in the temporal dimension, instead of using cross-attention between speech and motion [12], we directly concatenate the motion features with the speech embeddings as well as other optional temporal conditions along the channel dimension, resulting in a natural temporal alignment between audio and motion as well as an exemption on the use of attention masks in cross-attention temporal conditioning [12]. For feature fusion, instead of using simple linear layer [53] which leads to slow convergence and unstable training, we design the motion-speech fusion residual block, which fuses concatenated features and projects them into the same shape as motion features as shown in Figure 2. The block consists of a LayerNorm (LN) and an MLP, where the LN ensures numerical stability, and the MLP is designed to predict the residual of the original motion features to facilitate fast convergence (Section 12).
Style-aware Transformer Block. This block injects global conditions (style and diffusion step $t$) to the fused features from the previous Motion-Speech Block. Firstly, style (person ID in this paper) and diffusion step $t$ are projected to vectors of the same shape by MLPs. Then, the summation of two global condition vectors will be fed into AdaIN [19]. Two AdaIN stylization blocks are inserted after self-attention and MLP in a Transformer encoder block. The AdaIN blocks apply the feature statistics substitution according to the new statistics computed from the fused global condition vector. Note that we utilize linear self-attention from [36] as an alternative for full self-attention to save computation and facilitate the inference speed.
@@
Note that to prevent the gradient from the gesture branch from affecting the expression encoder, we cut off the gradient of the predicted expression $\hat{x}_{0(t)}^{E}$ before passing it to the gesture encoder, as illustrated in Figure 2. The Uni-EG module will repeat n times to get the final noise in each forward of the network.
Figure 3: Illustration for outpainting-based arbitrary long sequence inference. We adapt the Repaint algorithm to DDIM sampling and achieve real-time smooth motion generation. Each blue bar represents a motion clip from a single sampling process.
@@ 3.4 Training
Loss functions. Except for the noise prediction loss in Equation 4, we also introduce the velocity loss $\mathcal{L}_{v}$ and Huber motion reconstruction loss $\mathcal{L}_{\delta}$. Since the two losses are computed on the motion $\hat{\mathbf{x}}_{0}$, we first compute the predicted motion $\hat{\mathbf{x}}_{0(t)}$ from the predicted noise $\hat{\mathbf{\epsilon}_{t}}$, which is similar to Equation 5.
Then, we have velocity loss computed as the mean square error of the velocity:
$\mathcal{L}_{v}=\mathbb{E}\left[\left\|(\mathbf{x}_{0}[1:]-\mathbf{x}_{0}[:-1])-(\mathbf{x}_{0}^{\prime}[1:]-\mathbf{x}_{0}^{\prime}[:-1])\right\|^{2}\right],$ (6)
and the Huber loss for the reconstruction of motion
@@ 3.5 Arbitrary-long Motion Generation
Fast Outpainting-based Partial Autoregressive Sampling (FOPPAS). As illustrated in Figure 3, starting from the second clip, we can fix the initial frames to be the same as the last frames of the previous clip and then outpaint the remaining frames in this clip. This method conditions the current clip only on the part of the previous clip frames; therefore, we call it partial autoregressive sampling via outpainting. Unlike the RNN-based methods [26] or train-time autoregressive diffusion models [53, 46], the number of overlapping frames of FOPPAS can be flexibly set and changeable anytime instead of being required and fixed once trained. Those methods [26, 53, 46] also require an initial ”seed motion” to generate the first clip. However, it is not convenient for users to find such seed sequences or they just want to generate the first clip randomly. In contrast, we can generate the first clip with the overlapping number set as 0. In our experiment, we choose Repaint [29] to perform diffusion-based outpainting.
@@
Last-steps Refinement with Blending. Although applying outpainting already gets a very smooth transition between clips, we occasionally observe slight inconsistency at the boundary. To refine and facilitate the consistency at the clip boundary, we perform a linear blending at the overlapping part of clips at the last two sampling steps. For the details of our fast outpainting algorithm, please refer to Algorithm 1 in the Appendix.
Figure 4: Qualitative Results of DiffSHEG on the BEAT [26] Dataset. (a) Facial expressions and Lip Sync. When provided with speech audio, our model produces expressions accompanied by precise lip shapes. (b) Gestures. Our method generates expressive and lifelike gestures to effectively communicate concepts like “dream”, “design” and “street”. Ours also generates rhythmic synchronization and emphasis, exemplified by gestures corresponding to “never” and “some”. The characters are rendered using Unreal Engine and Metahuman. Please see the supplementary and demo video for more qualitative results.
Figure 5: Qualitative Comparison on BEAT [26] Dataset. In comparison to baseline methods, our approach generates a broader range of natural, agile, and diverse gestures that are closely synchronized with the audio input. When saying ”journalist”, the character driven by our motion raises double hands to stress this word; When saying ”never”, our motion shows two times up-and-down right hand and fingers, corresponding to the two syllables “ne” and “ver”.
@@ 4.2 Experimental Setup
Baselines. We compare our method with CaMN [26], DiffGesture [53], DiffuseStyleGesture [46] and LDA [1] on BEAT. CaMN is proposed with the BEAT dataset, which is based on LSTM and can fuse multiple conditions, including audio, text, face, and emotion. DiffGesture, DiffuseStyleGesture, and LDA are newly proposed diffusion-based co-speech generation methods. For gesture generation, the four baselines are retrained on the axis-angle rotation representation instead of the Euler angles. For expression generation, since four baseline methods were originally proposed for gesture generation solely, we substitute the gesture into the expression data to generate expressions independently. For a fair comparison, we condition all the models on audio and person ID. On the SHOW dataset, we utilize two baselines that focus on joint expression and gesture generation: TalkSHOW [48] and the method proposed in [14] named LS3DCG here. Since all the baselines only focus on the upper body movement for gesture generation, we follow this setting despite that our framework also works for the lower body motion.
Figure 8: Results of the user study. This shows user preference percentage in terms of four metrics: realism, gesture-speech synchronism, expression-speech synchronism, and motion diversity. In both datasets and all metrics, our method is dominantly preferred. DSG and DG are the abbreviations of DiffuseStyleGesture and DiffuseGesture.

### 06 Apr 2024 v2

@@ 1 Introduction
In this work, we propose DiffSHEG, a unified Diffusion-based Holistic Speech-driven Expression and Gesture generation framework, illustrated in Figure 1. To capture the joint distribution, DiffSHEG utilizes diffusion models [16] with a unified expression-gesture denoising network. As shown in Figure 2, our denoising network consists of two modules, an audio encoder, and a Transformer-based Uni-direction Expression-Gesture (UniEG) generator, which has a unidirectional flow from expression to gesture. The proposed framework ensures a natural temporal alignment between speech and motion, leveraging the relationship between expressions and gestures. Furthermore, we introduce a Fast Out-Painting-based Partial Autoregressive Sampling (FOPPAS) method to synthesize arbitrary long sequences efficiently. FOPPAS enables real-time streaming sequence generation without conditioning on previous frames during training, providing more flexibility and efficiency.
@@
Figure 2: DiffSHEG framework overview. Left: Audio Encoders and UniEG-Transformer Generator. Given an audio clip, we encode the audio into a low-level feature Mel-Spectrogram and a high-level HuBERT feature. An audio encoder learns a mid-level representation of speech. The audio features are concatenated with other optional temporal conditions and then fed into the UniEG Transformer Denoiser. The denoising block fuses the conditions with noisy motion at diffusion step t and feeds it into style-aware transformers to get the predicted noises. The uni-directional condition flow is enforced from expression to gesture for joint distribution learning. Right: The detailed architecture of style-aware Transformer encoder and motion-condition fusion residual block.
@@ 2 Related Work
Co-speech Expression Generation. Co-speech Expression Generation, also known as talking head/face generation, has been a topic of interest in the field. Early related methods primarily relied on rule-based procedural approaches [10, 31, 42, 46], which required significant manual effort despite offering explicit and flexible control. With the advancements in deep learning, data-driven methods [7, 28, 21, 41, 33, 9, 20, 53, 36, 29] have emerged, focusing on generating images that correspond to audio input. However, these methods often produce distorted pixels and inconsistent 3D results. On another front, some approaches [21, 35] propose animating 3D faces. More recently, FaceFormer [12] and CodeTalker [45] have utilized Transformer models to animate 3D facial vertices using designed attention maps for temporal alignment. While lip movement is closely tied to speech, most data-driven methods for expression animation design deterministic mappings, lacking diversity in eye movement and facial expressions. In contrast to these methods, our approach aims to leverage generative models to generate diverse facial expressions with precise lip movements driven by speech.
@@
Holistic Co-Speech Expression and Gesture Generation. There are some recent methods [14, 49] exploring the holistic joint generation of expression and gesture. Habibie et al. [14] first try to synthesize 3D facial and gesture motion at the same time. They propose a CNN-based framework that shares the same speech encoder and uses three decoders to generate distinct whole-body keypoints (i.e., for face, hand, body), followed by a discriminator to provide adversarial loss during training. However, this method learns deterministic mapping, which violates the nature of many-to-many mapping from speech to gesture motion. Yi et al. [49] use an encoder-decoder network with pre-trained Wav2Vec [4] audio feature for expression generation and VQ-VAE for gesture generation. Nevertheless, they generate the expressions and gestures separately, overlooking the joint distribution between them. Moreover, methods based on VQ-VAE are constrained by a finite set of motion tokens, potentially limiting their ability to generate diverse and agile motions.
@@ 3 Method
How to model the expression-gesture joint distribution can be key to this task. A naïve way is directly concatenating the two vectors together and passing them to deep networks simultaneously, such that the features of both can be shared by each other. However, we empirically find this does not work well and leads to sub-optimal results in our experiment. We hypothesize that the information flow from gesture to expression would interfere with the mapping from speech to expression. Intuitively, expressions can serve as cues for gestures since the gestures usually behave consistently with the emotions and speech inferred from expression [40] and lip movement [11]. On the contrary, gestures can hardly affect expressions, especially the lips, which have a strong and exclusive correlation with speech. Imagine when you express surprise, you may have widened eyes, opened mouth, and raised eyebrows, with frozen or startled gestures such as clasping hands and leaning backward. Therefore, we propose a unified framework based on diffusion models with uni-directional information flow from expression to gesture for unified co-speech expression and gesture generation to capture their joint distributions. At inference time, to generate arbitrary-long motion sequences, we propose an fast out-painting-based partial autoregressive sampling (FOPPAS) strategy, which generates partially overlapping clips one by one with a smooth motion at the intersection. Additionally, FOPPAS also features its real-time streaming inference ability. In the following sections, we will introduce the problem formulation and diffusion model [16], and then elaborate on the design of our proposed framework.
@@ 3.1 Problem Formulation
Given an arbitrary-long (streaming) audio, we aim to generate realistic, synchronous, and diverse gestures as well as expressions. To get training samples, we cut the aligned audio and motion sequences into clips with the same time duration using sliding windows. For each $N$-frame clip, we encode the corresponding audio clip to audio features $\mathbf{A}=[{\mathbf{a}_{1},\dots,\mathbf{a}_{N}}]$. Gesture clip $\mathbf{G}=[{\mathbf{g}_{1},\dots,\mathbf{g}_{N}}]$ is represented by joint rotations in axis-angle format, where $\mathbf{g}_{i}\in\mathbb{R}^{3J}$ and $J$ is the total joint number. Expression clip $\mathbf{E}=[{\mathbf{e}_{1},\dots,\mathbf{e}_{N}}]$ is represented by the blend shape weights where $\mathbf{e}_{i}\in\mathbb{R}^{C_{exp}}$ and $C_{exp}$ is the channel dimension of expression. The holistic motion sequence $\mathbf{M}=Concat(\mathbf{G},\mathbf{E})$, where each motion frame $\mathbf{m}_{i}\in\mathbb{R}^{3J+C_{exp}}$. The training objective of our framework is to reconstruct the motion clip $\mathbf{M}$ conditioned on audio clip $\mathbf{A}$, which is similar to the Multi-Input Multi-Output (MIMO) setting [34]. At inference time, we generate realistic, diverse, and well-aligned co-speech expressions and gestures with motion clips smoothly connected.
@@ 3.2 Preliminary: Diffusion Models
DiffSHEG uses diffusion models [16] for this task which consists of a diffusion and a denoising process. For explicitness, we substitute notations of expression, gesture, and motion clip $\mathbf{E},\mathbf{G},\mathbf{M}$ with $\mathbf{x}_{0}^{E},\mathbf{x}_{0}^{G},\mathbf{x}_{0}$. Given a motion clip distribution $p(\mathbf{x}_{0})\in\mathbb{R}^{N\times(3J+C_{exp})}$, our goal is to train a model parameterized by $\theta$ to approximate $p(\mathbf{x}_{0})$.
Diffusion Process. In the diffusion process, models progressively corrupts input data $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$ according to a predefined schedule $\beta_{t}\in(0,1)$, eventually turning data distribution into an isotropic Gaussian in $T$ steps. Each diffusion transition can be assumed as
@@
where the full diffusion process can be written as
@@
Denoising Process. In the denoising process, models learn to invert the diffusion procedure so that it can turn random noise into real data distribution at inference. The corresponding denoising process can be written as
@@
$\mathcal{L}_{t}$ $=\mathbb{E}_{\mathbf{x}_{0},\epsilon}\left[\left\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t\right)\right\|^{2}\right].$ (4)
@@ 3.3 DiffSHEG Framework
Motion-Speech Fusion Residual Block. In order to condition the motion on speech and align them in the temporal dimension, instead of using cross-attention between speech and motion [12], we directly concatenate the motion features with the speech embeddings as well as other optional temporal conditions along the channel dimension, resulting in a natural temporal alignment between audio and motion as well as an exemption on the use of attention masks in cross-attention temporal conditioning [12]. For feature fusion, instead of using simple linear layer [54] which leads to slow convergence and unstable training, we design the motion-speech fusion residual block, which fuses concatenated features and projects them into the same shape as motion features as shown in Figure 2. The block consists of a LayerNorm (LN) and an MLP, where the LN ensures numerical stability, and the MLP is designed to predict the residual of the original motion features to facilitate fast convergence (Section 10 in Appendix).
Style-aware Transformer Block. This block injects global conditions (style and diffusion step $t$) to the fused features from the previous Motion-Speech Block. Firstly, style (person ID in this paper) and diffusion step $t$ are projected to vectors of the same shape by MLPs. Then, the summation of two global condition vectors will be fed into AdaIN [19]. Two AdaIN stylization blocks are inserted after self-attention and MLP in a Transformer encoder block. The AdaIN blocks apply the feature statistics substitution according to the new statistics computed from the fused global condition vector. Note that we utilize linear self-attention [37, 52] as an alternative for full self-attention to save computation and facilitate fast inference.
@@
Note that to prevent the gradient of the gesture branch from affecting the expression encoder, we cut off the gradient of the predicted expression $\hat{x}_{0(t)}^{E}$ before passing it to the gesture encoder, as illustrated in Figure 2. The Uni-EG module will repeat $n$ times to get the final noise in each forward of the network.
Figure 3: Illustration for outpainting-based arbitrary long sequence inference. Given a previous clip, we generate current clip by outpainting the remaining frames (light blue) according to the overlaping frames (deep blue). Each row of blue bar represents a motion clip from a single sampling process.
@@ 3.4 Training
Loss functions. Except for the noise prediction loss in Equation 4, we also introduce the velocity loss $\mathcal{L}_{v}$ and Huber motion reconstruction loss $\mathcal{L}_{\delta}$. Since the two losses are computed on the motion $\hat{\mathbf{x}}_{0}$, we first compute the predicted motion $\hat{\mathbf{x}}_{0(t)}$ from the predicted noise $\hat{\mathbf{\epsilon}_{t}}$, which is similar to Equation 5. Then, we have velocity loss computed as the mean square error of the velocity:
$\mathcal{L}_{v}=\mathbb{E}\left[\left\|(\mathbf{x}_{0}[1:]-\mathbf{x}_{0}[:-1])-(\hat{\mathbf{x}}_{0}[1:]-\hat{\mathbf{x}}_{0}[:-1])\right\|^{2}\right],$ (6)
and the Huber loss for the reconstruction of motion:
@@ 3.5 Arbitrary-long Motion Generation
Fast Outpainting-based Partial Autoregressive Sampling (FOPPAS). As illustrated in Figure 3, starting from the second clip, we can fix the initial frames to be the same as the last frames of the previous clip and then outpaint the remaining frames in this clip. This method conditions the current clip only on the part of the previous clip frames; therefore, we call it partial autoregressive sampling via outpainting. Unlike the RNN-based methods [26] or train-time autoregressive diffusion models [54, 47], the number of overlapping frames of FOPPAS can be flexibly set and changeable anytime instead of being required and fixed once trained. Those methods [26, 54, 47] also require an initial "seed motion" to generate the first clip. However, it is not convenient for users to find such seed sequences or they just want to generate the first clip randomly. In contrast, we can generate the first clip with the overlapping number set as 0. In our experiment, we choose Repaint [30] to perform diffusion-based outpainting.
@@
Last-steps Refinement with Blending. Although applying outpainting already gets a very smooth transition between clips, we occasionally observe slight inconsistency at the boundary. To refine and facilitate the consistency at the clip boundary, we perform a linear blending at the overlapping part of clips at the last two sampling steps. For the details of FOPPAS, please refer to Algorithm 1 in the Appendix.
Figure 4: Qualitative Comparison on BEAT [26] Dataset. In comparison to baseline methods, our approach generates a broader range of natural, agile, and diverse gestures that are closely synchronized with the audio input. When saying "journalist", the character driven by our motion raises double hands to stress this word; When saying "never", our motion shows two times up-and-down right hand and fingers, corresponding to the two syllables “ne" and “ver". The character is from MetaHuman [55] rendered by Unreal Engine 5 [56].
@@ 4.2 Experimental Setup
Baselines. We compare our method with CaMN [26], DiffGesture [54], DiffuseStyleGesture [47] and ListentDenoiseAction (LDA) [1] on BEAT. CaMN is proposed with the BEAT dataset, which is based on LSTM and can fuse multiple conditions, including audio, text, face, and emotion. DiffGesture, DiffuseStyleGesture, and LDA are newly proposed diffusion-based co-speech generation methods. For gesture generation, the four baselines are retrained on the axis-angle rotation representation instead of the Euler angles. For expression generation, since four baseline methods were originally proposed for gesture generation solely, we substitute the gesture into the expression data to generate expressions independently. For a fair comparison, we condition all the models on audio and person ID. On the SHOW dataset, we utilize two baselines that focus on joint expression and gesture generation: TalkSHOW [49] and the method proposed in [14] named LS3DCG here. Since all the baselines only focus on the upper body movement for gesture generation, we follow this setting despite that our framework also works for the lower body motion.
Figure 7: Results of the user study. The chart shows user preference percentage in terms of four metrics: realism, gesture-speech synchronism, expression-speech synchronism, and motion diversity. In both datasets and all metrics, our method is dominantly preferred. DSG and DG are the abbreviations of DiffuseStyleGesture and DiffuseGesture.